# Near-inertial echoes of ageostrophic instability in submesoscale filaments

### v1

@@ Abstract
Ocean submesoscales, flows with characteristic size around 10m–10km, are transitional between the larger, rotationally-constrained mesoscale and three-dimensional turbulence. In this paper we present simulations of a submesoscale ocean filament. The strong vertical transport associated with submesoscale fronts and filaments ventilates the deeper water and provides a route for the observed forward cascade of energy in the ocean. In our case, the filament is strongly sheared in both vertical and cross-filament directions and is unstable. Instability indeed dominates the early behaviour with a fast extraction of kinetic energy from the vertically sheared thermal wind. The filament is expected to be subject to instabilities, including symmetric and Kelvin-Helmholtz instabilities. An instability emerges with a circulation that does not match either, though symmetric modes become apparent later. The action of the instability is sufficiently rapid that the filament does not respond in a geostrophically balanced sense. Instead, it later exhibits vertically sheared near-inertial oscillations with higher amplitude as the initial minimum Richardson number decreases. Horizontal gradients strengthen only briefly as the fronts restratify. These unstable filaments can be generated by strong mixing events at pre-existing stable structures. The oscillations must also be considered when designing initial conditions for numerical experiments of submesoscale flows, where the initial state is intended to be balanced.
@@ 1 Introduction
The submesoscale ocean is dominated by highly horizontally anisotropic features. Fronts — confined lateral changes in density — remain balanced by the Coriolis acceleration of an along-front jet. In the open ocean, these features are created primarily at the edges of mesoscale vortices such as those created by western boundary currents. As they are drawn out by the strain flow in these regions, horizontal gradients in velocities and tracers become larger, eventually leading to the onset of a new dynamical regime, where geostrophic and hydrostatic balances no longer describe the flow well (Taylor & Thompson, 2022; Barkan et al., 2019).
@@
The strain flow and turbulence drive an ageostrophic secondary circulation around the front (McWilliams, 2021; Sullivan & McWilliams, 2018). The resulting vertical velocities provide a route for transport of tracers, momentum and energy between the surface layer and deep water (Klein et al., 2019; Verma et al., 2019). Particularly, the submesoscale is a link between the larger scale, horizontally non-divergent flows and smaller scale turbulence that facilitates the observed forward cascade of energy in the ocean (Srinivasan et al., 2023).
@@
In the absence of turbulence, horizontal gradients will approach infinity in finite time when forced by a strain flow (Hoskins & Bretherton, 1972). For a realistic front, some effective horizontal diffusion due to turbulence will arrest the frontogenesis before this singularity is reached, though this process is not well-understood. The turbulence, namely the vertical momentum mixing induced by surface cooling or wind stress, acts on the vertically sheared thermal wind velocity. If only these processes act in the submesoscale mixed layer, the expected momentum balance is the turbulent thermal wind balance between the Coriolis acceleration of an along-front jet, the horizontal pressure gradient and vertical momentum mixing (Gula et al., 2014; Bodner et al., 2023).
This study investigates the evolution of a dense filament resembling the configuration of Sullivan & McWilliams (2018), in the initially low-turbulence regime. A filament is a pair of opposing, interacting fronts. The vertical vorticity of a dense filament in thermal wind balance is large and positive in the centre between two smaller magnitude negative vorticity regions.
@@
Of particular interest in this study is symmetric instability (SI). A symmetrically unstable flow is unstable to perturbations that exchange fluid parcels along isopycnals, with no variation (hence symmetric; Stone, 1966) along the direction of mean flow. The instability manifests as highly anisotropic rolls, along the direction of the frontal jet, with major axis aligned with isopycnals. SI extracts kinetic energy from the thermal wind velocity shear (Taylor & Thompson, 2022; Haine & Marshall, 1998).
@@
For a Boussinesq fluid with equal momentum and buoyancy diffusivities, the stability of a flow to symmetric perturbations can be diagnosed by the sign of Ertel potential vorticity $q$. Fluid unstable to SI has $fq=f(f\hat{\boldsymbol{z}}+\nabla\times\boldsymbol{u})\cdot\nabla b<0,$ where $\boldsymbol{u}=u\hat{\boldsymbol{x}}+v\hat{\boldsymbol{y}}+w\hat{\boldsymbol{z}}$ and $b$ are velocity and buoyancy fields respectively and $f$ is the Coriolis frequency (Emanuel, 1979). SI competes with gravitational, inertial, baroclinic and Kelvin-Helmholtz instabilities, and dominates in regions where the Richardson number of the flow lies in the range $0.950.25<Ri<$ 0.95 (Stone, 1966; Thomas et al., 2013). The small size and strong baroclinicity of submesoscale fronts often satisfies the condition for SI (Thomas et al., 2013; Gula et al., 2014).
@@
In this paper, we present a set of simulations, unstable to various instabilities owing to the low Richardson number ($Ri\leq 0.2$). Our results focus on the spin-down of the fronts and their subsequent instability. We show that instabilities act quickly to remove kinetic energy and excite near-inertial oscillations within the fronts that briefly strengthen horizontal buoyancy gradients at the surface. Finally, we argue that numerical studies of similar features must be designed with this instability and subsequent oscillations in mind, or risk drawing conclusions that may not be generalisable to balanced flows in the real ocean.
2 Model
@@
$\frac{\text{D}\boldsymbol{u}}{\text{D}t}+f\boldsymbol{\hat{z}}\times\boldsymbol{u}=-\nabla\phi+b\boldsymbol{\hat{z}}+\nabla\boldsymbol{\cdot}(\nu\nabla\boldsymbol{u})+\boldsymbol{F}_{\boldsymbol{u}},$ ($)1𝑎$
$\frac{\text{D}b}{\text{D}t}=\nabla\boldsymbol{\cdot}(\kappa\nabla b)+F_{b},\quad\text{and}\quad\nabla\boldsymbol{\cdot}\boldsymbol{u}=0,$ ($)1𝑏𝑐$
with geopotential $\phi$, kinematic eddy viscosity $\nu$, and buoyancy diffusivity $\kappa$. The terms $\boldsymbol{F_{u}}$ and $F_{b}$ are velocity and buoyancy forcing terms. We elaborate on these quantities below. The equation set (($)1𝑎$) is integrated using Oceananigans, a Julia package for simulation of incompressible fluid flows in an oceanic context (Ramadhan et al., 2020).
Figure 1: a) Schematic of the simulation setup and an example initial condition (not to scale). Black contours are isopycnals and the heatmap shows the form of the balanced thermal wind velocity. The vertical grid size smoothly transitions between two uniform values, such that about 96 grid cells are in the mixed layer, and 32 grid cells comprise the remainder of the vertical domain (not shown). b) The local Richardson number for the reference state with $Ri_{\text{min}}=0.1$. Fluid inside the red, dashed contour has $fq<0$.
A schematic of the simulation setup is shown in figure 1, and details of the initial condition are in Appendix A. The grid size is $(N_{x},N_{y},N_{z})=(1024,1024,128)$ and the simulation domain is $(x,y)\in[-5L,5L)^{2}$, where $L$ is a measure of the initial surface separation between the centres of the two counter-flowing jets, and $z\in[-2.5H,0]$, where $H$ is the initial depth of the mixed layer. The grid spacing is uniform in the horizontal directions. The vertical grid spacing smoothly transitions between two uniform values at the pycnocline, such that approximately $3/4$ of the cells are in the surface mixed layer ($z>-H$), while the remainder are in the deep water. Quantities are non-dimensionalised such that $f=L=1$, although we sometimes keep the notations $f$ and $L$ for physical clarity.
The domain is periodic in both horizontal directions. In the vertical, boundary conditions are no-flux — the filament is not forced by wind stress or surface cooling, namely,
$u_{z}=v_{z}=w=b_{z}=0\quad\text{at}\quad z=\{-2.5H,0\},$
where $\phi_{i}$ denotes differentiation of $\phi$ with respect to coordinate $i$.
@@
The isotropic eddy viscosity $\nu$ is parameterised as a Smagorinsky-Lilly turbulence closure with a Smagorinsky constant $C=0.16$ (Smagorinsky, 1963; Lilly, 1967), and we set $\kappa=\nu$. The adaptive time-stepping routine aims to achieve a maximum Courant-Friedrichs-Levy number of 0.15.
3 Experiments
@@
The reference buoyancy contains a weakly-stratified surface layer above a strongly-stratified fluid. The height of the transition between these two stratifications varies, providing the horizontal buoyancy gradient that makes up the filament. The exact definition of the functional forms of $v_{0}$ and $b_{0}$ are given in appendix A.
@@
$\overline{\phi}(x,z,t):=\frac{1}{10L}\int_{-5L}^{5L}\phi(x,y,z,t)\;\text{d}y.$ (4)
@@
The primary non-dimensional numbers that determine the dynamical regime of a submesoscale mixed layer flow are the Rossby, Richardson and (turbulent) Ekman numbers (Bodner et al., 2023). For the base state, the local values of each are, respectively,
$Ro=\frac{1}{f}\frac{\partial v_{0}}{\partial x},\quad Ri=\frac{\partial b_{0}/\partial z}{(\partial v_{0}/\partial z)^{2}}\quad\text{and}\quad Ek=\frac{\partial\overline{w^{\prime}v^{\prime}}/\partial z}{v_{0}f}.$ (5)
@@
With no surface forcing, the filament is initially in a low-turbulence ($Ek\sim 0$) regime. We parametrise our initial conditions by extremal values of the local Rossby and Richardson numbers as follows.
@@
This study is concerned with the submesoscale ocean, where advection and the Coriolis acceleration are comparable in magnitude: $Ro\sim 1$. To limit the number of possible instabilities, the simulations have minimum vertical vorticity above $-f$, which excludes inertial instability. We use $Ro_{\text{min}}=-0.8$ here. Given the prescribed filament shape, this results in a maximum $Ro_{\text{max}}=1.9$. For typical horizontal velocities in the ocean this corresponds to a jet width of order $1-10~{}\text{km}$ at mid-latitudes.
@@
The submesoscale mixed layer often has $Ri\lesssim 1$. For the filament shape used here, the Richardson number is approximately constant with depth in the mixed layer, and attains a minimum value in the centre of each front. We present the results of a set of low-Ri filaments, with a range of minimum Richardson numbers: $Ri_{\text{min}}\in\{0.0,0.1,0.2\}$.
The remainder of the non-dimensional parameters, as well as the relationship between $Ro_{\text{min}}$, $Ri_{\text{min}}$ and dimensional parameters, is given in appendix A.
Figure 1b shows the local $Ri$ of the reference state, and the region with $fq<0$ at the outer edges of the filament.
Before we impose the reference state, there is a brief initialisation process. We pre-initialise the simulation at $t=-1$ with a state identical to the reference state at $|x|\rightarrow\infty$. We then add noise to all velocities and run the simulation until $t=0$. This produces fields $\boldsymbol{u}_{\text{init}}=\overline{\boldsymbol{u}}_{\text{init}}+\boldsymbol{u}_{\text{init}}^{\prime}$ and $b_{\text{init}}=\overline{b}_{\text{init}}+b_{\text{init}}^{\prime}$. The initial state of the filament phase of the simulation is taken to be the reference state plus the fluctuation terms from the initialisation phase, namely,
@@
The mean, balanced state provides the energy source for instabilities with the nature of the instability determining the dominant energy source. Energy transfer from fluctuations to the mean kinetic energy of a laterally or vertically sheared velocity profile is given by the lateral or vertical shear production respectively
$\text{LSP}=\overline{\boldsymbol{u}}_{x}\boldsymbol{\cdot}\overline{u^{\prime}\boldsymbol{u}^{\prime}},\quad\text{VSP}=\overline{\boldsymbol{u}}_{z}\boldsymbol{\cdot}\overline{w^{\prime}\boldsymbol{u}^{\prime}}.$ (6)
Mean potential energy is transferred via the buoyancy flux,
$\text{BFLUX}=-\overline{w^{\prime}b^{\prime}}.$ (7)

### v2

@@ Abstract
Ocean submesoscales, flows with characteristic size around $10~{}\text{m}-10~{}\text{km}$, are transitional between the larger, rotationally-constrained mesoscale and three-dimensional turbulence. In this paper we present simulations of a submesoscale ocean filament. In our case, the filament is strongly sheared in both vertical and cross-filament directions and is unstable. Instability indeed dominates the early behaviour with a fast extraction of kinetic energy from the vertically sheared thermal wind. The filament is expected to be subject to instabilities, including symmetric and Kelvin-Helmholtz instabilities. However, the instability that emerges does not exhibit characteristics that match either, and appears to be non-normal in nature. The action of the instability is sufficiently rapid — with energy extraction from the mean flow emerging and peaking within the first inertial period ($\sim 18~{}\text{hr}$) — that the filament does not respond in a geostrophically balanced sense. Instead, it later exhibits vertically sheared near-inertial oscillations with higher amplitude as the initial minimum Richardson number decreases and at all initial noise levels. Horizontal gradients strengthen only briefly as the fronts restratify. These unstable filaments can be generated by strong mixing events at pre-existing stable structures; we also caution against inadvertently triggering this response in idealised studies that start in a very unstable state.
@@ 1 Introduction
At length scales larger than $\sim 100\text{ km}$, flows in the ocean are close to geostrophic balance — a state of approximate balance between the Coriolis acceleration and pressure gradient — and evolve slowly, over many days. At smaller length scales, beyond those that can currently be resolved in global ocean models, lies the submesoscale ($\sim 0.1-10\text{ km}$). Understanding of the impact of submesoscale flows on the ocean is emerging, with observational and numerical evidence that processes at this scale, which is poorly represented in parametrisations, are crucial components of the ocean’s physical and biogeochemical processes (Mahadevan, 2016; Calil et al., 2021; Gula et al., 2022; Aravind et al., 2023; Pham et al., 2024; Zhu et al., 2024).
@@
Images of ocean properties at submesoscale resolution are often peppered with highly anisotropic features. Fronts — confined lateral gradients in density — and filaments — opposing pairs of fronts — remain balanced by the Coriolis acceleration of an along-front jet. In the open ocean, these features are created primarily at the edges of mesoscale vortices such as those created by separating western boundary currents, such as the Gulf Stream or Kuroshio current (McWilliams, 2016). As they are drawn out by the strain flow in these regions, horizontal gradients in velocities and tracers become larger, eventually leading to the onset of a new dynamical regime, where geostrophic and hydrostatic balances no longer describe the flow well (Shakespeare, 2016; Barkan et al., 2019; Taylor & Thompson, 2022).
@@
In addition to the along-front geostrophic jet, an ageostrophic secondary circulation may develop at a front due to a background strain, turbulent mixing, instability or otherwise (Sullivan & McWilliams, 2018; Crowe & Taylor, 2018; McWilliams, 2021). The resulting vertical velocities provide a route for transport of tracers, momentum and energy between the surface layer and deep water (Mahadevan, 2016; Klein et al., 2019; Verma et al., 2019; Freilich & Mahadevan, 2021). Particularly, the submesoscale is a link between the larger scale, horizontally non-divergent flows and smaller scale turbulence that facilitates the observed forward cascade of energy in the ocean (Srinivasan et al., 2023).
@@
In the absence of turbulence, horizontal gradients will approach infinity in finite time when forced by a strain flow (Hoskins & Bretherton, 1972). For a realistic front, some effective horizontal diffusion due to turbulence will arrest the frontogenesis before this singularity is reached, though this process is not well-understood (Samelson & Skyllingstad, 2016; Verma et al., 2019). The turbulence, namely the vertical momentum mixing induced by surface cooling or wind stress, acts on the vertically sheared thermal wind velocity. If only these processes act in the submesoscale mixed layer, the expected momentum balance is the turbulent thermal wind balance between the Coriolis acceleration of an along-front jet, the horizontal pressure gradient and vertical momentum mixing (Gula et al., 2014; Crowe & Taylor, 2018; Bodner et al., 2023).
@@
Of particular interest in this study is symmetric instability (SI). A symmetrically unstable flow is unstable to perturbations that exchange fluid parcels along isopycnals, with no variation along the direction of mean flow (hence symmetric; Stone, 1966). The instability manifests as highly anisotropic rolls, along the direction of the frontal jet, with major axis aligned with isopycnals. The most unstable across-front wavenumber tends to infinity in the inviscid limit, so the wavenumber of the observed mode is presumed to be restricted by diffusion (Harris et al., 2022). SI extracts kinetic energy from the thermal wind velocity shear (Haine & Marshall, 1998; Thomas et al., 2013; Taylor & Thompson, 2022).
@@
For a Boussinesq fluid with equal momentum and buoyancy diffusivities, the stability of a flow to symmetric perturbations can be diagnosed by the sign of Ertel potential vorticity $q$. Fluid unstable to SI has $fq=f(f\hat{\boldsymbol{z}}+\nabla\times\boldsymbol{u})\cdot\nabla b<0,$ where $\boldsymbol{u}=u\hat{\boldsymbol{x}}+v\hat{\boldsymbol{y}}+w\hat{\boldsymbol{z}}$ and $b$ are velocity and buoyancy fields respectively and $f$ is the Coriolis frequency (Emanuel, 1979). SI competes with gravitational, inertial, baroclinic and Kelvin-Helmholtz instabilities, and dominates in regions where the Richardson number ($Ri=N^{2}/S^{2}$ where $N$ and $S$ are buoyancy frequency and vertical shear of horizontal velocity respectively) of the flow lies in the range $0.950.25<Ri<$ 0.95 (Stone, 1966; Thomas et al., 2013). The small size and strong baroclinicity of submesoscale fronts often produces favourable conditions for SI (Thomas et al., 2013; Gula et al., 2014).
The behaviour of fluid susceptible to SI (or close to unstable) is suspected to have important consequences for ocean flows at the submesoscale and, for example, has been studied in the contexts of frontogenesis and frontal arrest (Thomas, 2012; Verma et al., 2019), mixing and vertical transport (Brannigan, 2016; Bosse et al., 2021), and near-inertial wave generation (Wienkers et al., 2021). However, the initial and long-time behaviour of a symmetrically unstable flow differ in form, growth rate and energy source (Heifetz & Farrell, 2007; Zemskova et al., 2020; Kimura, 2024). Initial motions in an unstable flow seeded with noise do not resemble the symmetric normal modes, and exhibit a faster growth rate across parameter space, a behaviour that is general across hydrodynamic stability (Trefethen et al., 1993). This non-normal, transient growth will play a particularly important role in this study.
@@
In this paper, we present a set of strongly baroclinic ($Ri\leq 0.2$) simulations resembling the configuration of Sullivan & McWilliams (2018), but with coarser spatial resolution and without the surface cooling that drives vertical mixing. Our results focus on the spin-down of the fronts in the presence of noise and their subsequent instability. We show that instabilities act quickly to remove kinetic energy and excite near-inertial oscillations within the fronts that briefly strengthen horizontal buoyancy gradients at the surface. We argue that we observe a transient response of the filament associated with its symmetric instability. Finally, we suggest that initial conditions of numerical studies of similar strongly baroclinic features must be designed with initial instability and possible subsequent oscillations in mind, or risk drawing conclusions that may not be generalisable to balanced flows in the real ocean.
2 Numerical simulations
@@
$\frac{\text{D}\boldsymbol{u}}{\text{D}t}+f\boldsymbol{\hat{z}}\times\boldsymbol{u}=-\nabla\varphi+b\boldsymbol{\hat{z}}+\nabla\cdot(\nu\nabla\boldsymbol{u})+\boldsymbol{F}_{\boldsymbol{u}},$ ($)1𝑎$
$\frac{\text{D}b}{\text{D}t}=\nabla\cdot(\nu\nabla b)+F_{b},\quad\text{and}\quad\nabla\boldsymbol{\cdot}\boldsymbol{u}=0,$ ($)1𝑏𝑐$
with $f=10^{-4}~{}\text{s}^{-1}$, geopotential $\varphi$ and Lagrangian derivative ${\text{D}\phi}/{\text{D}t}=\partial\phi/\partial t+\boldsymbol{u}\cdot\nabla\phi$. The terms $\boldsymbol{F_{u}}$ and $F_{b}$ are velocity and buoyancy forcing terms. We elaborate on these quantities below. The equation set (($)1𝑎$) is integrated using Oceananigans, a Julia package for simulation of incompressible fluid flows in an oceanic context (Ramadhan et al., 2020).
Figure 1: a) The local Rossby number for the reference state with $Ri_{\min{}}=0.1$. Black contours are isopycnals. Fluid inside the red, dashed contour has $fq<0$. b) The local Richardson number, as in a).
The grid size is $(N_{x},N_{y},N_{z})=(1024,1024,128)$ and the simulation domain is $(x,y)\in[-5L,5L)^{2}$, where $L=1000\text{ m}$ is a measure of the initial surface separation between the centres of the two counter-flowing jets, and $z\in[-2.5H,0]$, where $H=100\text{ m}$ is the initial depth of the mixed layer. The grid spacing is uniform in the horizontal directions. The vertical grid spacing smoothly transitions between two uniform values at the pycnocline, such that approximately $3/4$ of the cells are in the surface mixed layer ($z>-H$), while the remainder are in the deep water.
The domain is periodic in both horizontal directions. In the vertical, boundary conditions are no-flux — the filament is not forced by wind stress or surface cooling, namely, $u_{,z}=v_{,z}=w=b_{,z}=0$ at $z=0$ and $u_{,z}=v_{,z}=w=0$ and $b_{,z}=N_{0}^{2}$ at $z=-2.5H$, where $\phi_{,i}$ denotes differentiation of $\phi$ with respect to coordinate $i$ and $N_{0}$ is the deep water buoyancy frequency.
@@
The isotropic eddy viscosity $\nu$ is parametrised as a Smagorinksy-Lilly turbulent closure with a Smagorinsky constant $C=0.16$ (Smagorinsky, 1963; Lilly, 1967), and we set $\kappa=\nu$. The adaptive time-stepping routine aims to achieve a maximum Courant-Friedrichs-Levy number of 0.5.
@@
The reference buoyancy contains a weakly-stratified surface layer above a strongly-stratified fluid. The height of the transition between these two stratifications varies, providing the horizontal buoyancy gradient that makes up the filament. The reference filament state is shown in figure 1, and the exact definition of the functional forms of $v_{0}$ and $b_{0}$ are given in appendix A.
@@
$\overline{\phi}(x,z,t):=\frac{1}{10L}\int_{-5L}^{5L}\phi(x,y,z,t)\;\text{d}y,$ (4)
@@
and down-front fluctuations $\phi^{\prime}=\phi-\overline{\phi}$. The primary non-dimensional numbers that determine the dynamical regime of an inviscid submesoscale mixed layer flow are the Rossby and Richardson numbers (McWilliams, 2016; Bodner et al., 2023). For the base state, the local values of each are, respectively,
$Ro=\frac{1}{f}\frac{\partial v_{0}}{\partial x},\quad\text{and}\quad Ri=\frac{\partial b_{0}/\partial z}{(\partial v_{0}/\partial z)^{2}}.$ (5)
@@
We parametrise our initial conditions by extremal values of the local Rossby and Richardson numbers as follows.
@@
This study is concerned with the submesoscale ocean, where advection and the Coriolis acceleration are comparable in magnitude: $Ro\sim 1$. To limit the number of possible instabilities, the reference state has minimum vertical vorticity above $-f$, which excludes inertial instability. We use $Ro_{\text{min}}=-0.8$ here. Given the prescribed filament shape, this results in a maximum $Ro_{\text{max}}=1.9$. For typical horizontal velocities in submesoscale fronts this corresponds to jet widths of order $1-10~{}\text{km}$ at mid-latitudes (Siegelman et al., 2020; Calil et al., 2021).
@@
The submesoscale mixed layer often has $Ri\lesssim 1$ (Zhu et al., 2024). For the filament shape used here, the Richardson number is approximately constant with depth in the mixed layer, and attains a minimum value in the horizontal centre of each front. We present the results of a set of low-Ri filaments, with a range of minimum Richardson numbers: $Ri_{\text{min}}\in\{0.0,0.1,0.2\}$. The remainder of the non-dimensional parameters, as well as the relationship between $Ro_{\text{min}}$, $Ri_{\text{min}}$ and dimensional parameters, is given in appendix A. Figures 1a and 1b show the local $Ro$ and $Ri$ of the $Ri_{\min{}}=0.1$ reference state respectively, and the region with $fq<0$ at the outer edges of the filament.
Figure 2: a) A slice of vertical velocity $w_{\text{init}}$ in the initial condition, just before the filament is imposed at $t=0$. b) Kinetic energy $\boldsymbol{u}_{\text{init}}\boldsymbol{\cdot}\boldsymbol{u}_{\text{init}}/2$ at the same time, horizontally averaged, as a function of depth.
Before we impose the reference state, there is a brief initialisation process to ensure that the initial conditions, with noise, satisfy the Boussinesq equations to at least first order, as well as to allow the sub-grid parametrisation to reduce very large gradients caused by the pre-initial random noise before the unstable flow is introduced. We pre-initialise the simulation at $ft/2\pi=-1$ with a state identical to the reference state at $|x|\rightarrow\infty$. We then add noise to all velocities and run the simulation until $t=0$. This produces fields $\boldsymbol{u}_{\text{init}}=\overline{\boldsymbol{u}}_{\text{init}}+\boldsymbol{u}_{\text{init}}^{\prime}$ and $b_{\text{init}}=\overline{b}_{\text{init}}+b_{\text{init}}^{\prime}$. The pre-initialised noise is sampled from a Gaussian distribution and is chosen to have similar magnitude and variation with depth of the turbulent kinetic energy to that which would be produced by forcing due to surface cooling of order $\sim 100\text{ Wm}^{-2}$ (see Sullivan & McWilliams, 2018). A slice of vertical velocity and the horizontally-averaged kinetic energy is shown in figure 2. We explore the impact of the choice of initial noise on the conclusions of this study in appendix B The initial state of the filament phase of the simulation is taken to be the reference state plus the fluctuation terms from the initialisation phase, namely,
@@
The mean, balanced state provides the energy source for instabilities with the nature of the instability determining the dominant energy source. We consider contributions to the rate of change of the total (kinetic and potential) energy in the mean state, integrated over the whole domain $\mathcal{V}$. For the boundary conditions considered here, the rate of change of mean total energy may be partitioned into turbulent flux terms as follows,
$\frac{\text{d}}{\text{d}t}\int_{\mathcal{V}}\text{d}V\left(\frac{1}{2}\overline{\boldsymbol{u}}\boldsymbol{\cdot}\overline{\boldsymbol{u}}-\overline{b}z\right)\approx\int_{\mathcal{V}}\text{d}V\bigg{(}{\overline{\boldsymbol{u}}_{,x}\boldsymbol{\cdot}\overline{u^{\prime}\boldsymbol{u}^{\prime}}}+{\overline{\boldsymbol{u}}_{,z}\boldsymbol{\cdot}\overline{w^{\prime}\boldsymbol{u}^{\prime}}}-{\overline{w^{\prime}b^{\prime}}}\bigg{)},$ (6)
wherein the direct impact of the sponge forcing and sub-grid dissipation on the reference state are assumed to be negligible. Energisation of fluctuations by a laterally or vertically sheared velocity profile is given by the lateral or vertical shear production respectively
$\text{LSP}=-\int_{\mathcal{V}}\text{d}V\;\overline{\boldsymbol{u}}_{,x}\boldsymbol{\cdot}\overline{u^{\prime}\boldsymbol{u}^{\prime}},\quad\text{VSP}=-\int_{\mathcal{V}}\text{d}V\;\overline{\boldsymbol{u}}_{,z}\boldsymbol{\cdot}\overline{w^{\prime}\boldsymbol{u}^{\prime}}.$
The VSP approximates the geostrophic shear production (GSP) if the down-front averaging can be assumed to separate the geostrophic and ageostrophic components. GSP provides the energy source for symmetric instability, among others. Lateral shear is known to energise inertial and barotropic instabilities. Mean potential energy is transferred to fluctuations via the (vertical) buoyancy flux, typically associated with baroclinic and gravitational instabilities,
$\text{BFLUX}=\int_{\mathcal{V}}\text{d}V\;\overline{w^{\prime}b^{\prime}}.$